Training Data

Positive Training (22 cases)
avalanche  bomb threat  broken leg  burglary  car collision
death threat  fire  flood  gas leak  heart attack
hurricane  landslide  murder  overdose  pneumonia
rape  roof collapse  sinking ship  stroke  tornado
train wreck  trapped miners
Negative Training (25 cases)
arthritis  broken dishwasher broken toe  cat in tree  contempt of court
dandruff  delayed train  dizziness  drunkenness enumeration
flat tire  frog  headache  leaky faucet littering
missing dog  paper cut  practical joke rain  roof leak
sore throat  sunset  truancy  vagrancy  vulgarity
Anchors  (6 dimensions)
crime happy help safe urgent

wash

Testing Results

Positive tests  Negative tests
Positive assault, coma, menopause, prank call,
Predictions  electrocution, heat stroke, pregnancy, traffic jam

homicide, looting, meningitis, robbery, suicide

Negative sprained ankle acne, annoying sister,
Predictions  campfire, desk,

mayday, meal

Accuracy  15/20 = 75.00%